Hello,
I have a list of equations below theres suppose to be two sets of parallel lines which one are ? also there are suppose to be 3 pairs of perpindicular lines as well i'm suppose to identify the perpindicular equations by letters and equations and for the parallel i'm suppose to criss cross by connecting each parallel line: heres the list::
correct me if i'm wrong
x = 3
y = 2x-1
y/2 = x+3
y = -7

2y = -x
x = -2

I know -2 and -7 are horizontial x=3 i think is perpindicular and 2x-1 is parallel so how can answer this problem logically to supply the answers that are needed thanks for looking and trying to help!

Answer by mmm4444bot(95)

Hello There:
Here is the information that you need to answer these questions.
An equation of the form x = c, where c is a real number, has a graph that's a vertical line.
Your two equations x = -2 and x = 3 are vertical lines that intersect the x-axis at (-2, 0) and (3, 0).
An equation of the form y = c, where c is a real number, has a graph that's a horizontal line.
Your equation y = -7 is a horizontal line that intersects the y-axis at (0, -7)
When two lines have the same slope, they are parallel.
When two lines have slopes that are negative reciprocals of each other, they aer perpendicular.
Solving for y shows us the slope.
This equation is already solved for y:
y = 2*x - 1
Solving this equation: y/2 = x + 3 gives us:
y = 2*x + 6
Both of these lines have slope 2, so they are parallel.
Solving 2*y = -x leads to: y = -(1/2)*x
This slope is -1/2; so it's the negative reciprocal of 2 (the slope of the parallel lines).
So, 2*y = -x is the line that is perpendicular to y/2 = x + 3 and y = 2*x - 1.
~ Mark